Find the equation to the straight line cutting of an intercept (-3) from the Y axis and inclined at an angle tan-1 3/5 to the X axis

Grade:10

1 Answers

Aditya Gupta
4 years ago
the eqn of a st line is 
y= mx+c where m= tantheta is the slope. since angle of inclination to x axis is given to be theta= tan-1 3/5 , so the slope m= tantheta= 3/5.
now, c is the y intercept.
so eqn: y= 3x/5 – 3
or 5(y+3)= 3x
kindly approve :)